# OpenRT: An Open-Source Red Teaming Framework for Multimodal LLMs

### v1

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) qwen3max; gpt5-2; ClaudeHaiku45; Gemini3ProPreview; comanici2025gemini; Meta2025Llama4; Grok41Fast2025 are increasingly powering real-world applications, including conversational assistants akhoroz2025conversational; fischer2024grillbot, code copilots ross2023programmer, and search agents gong2024cosearchagent; xi2025survey. To mitigate harmful behavior, these systems are often equipped with safety alignment mechanisms lab2025safework; qi2024safety; wang2025safevid and safeguard policies inan2023llama; wang2024advqdet; wang2025tapt; zhao2025qwen3guard. Despite being widely adopted, conventional defenses such as system prompts sharma2025sysformer, safety filters huang2024vaccine; gong2025safety, and refusal-aware fine-tuning li2024self; zhang2025safety, remain vulnerable to adversarial attacks chen2025evolve; rahman2025x; wang2025safeevalagent, revealing a significant gap between perceived safety and empirical worst-case vulnerabilities.
@@
Despite significant advances in jailbreak techniques qi2024visual; zou2023universal; chao2023jailbreaking, the ecosystem for systematically evaluating adversarial robustness remains fragmented. Most existing red-teaming frameworks zhou2024easyjailbreak; chao2024jailbreakbench; mazeika2024harmbench; xu2024bag; jia2025omnisafebench focus on a narrow subset of attacks, limited threat models, or a small selection of target models. As the number and variety of red-teaming approaches grow, including evolutionary strategies chen2025evolve, multi-modal jailbreaks wang2025ideator; gong2025figstep, multi-turn optimization samvelyan2024rainbow; liu2024autodan, and multi-agent coordination rahman2025x; ren2024derail, the lack of a unified experimental framework has become a critical bottleneck. Such fragmentation undermines reproducible benchmarking and limits the systematic evaluation of vulnerabilities across models. As a result, it remains difficult to establish standardized baselines for attack efficacy or quantify the consistency of safety failures across diverse models.
@@
In this work, we introduce OpenRT, a modular and extensible framework designed for the red teaming of MLLMs. Unlike existing toolboxes often limited to a narrow subset of classic attacks, OpenRT supports massively parallel jailbreaking in both white-box and black-box settings, engineered specifically for high-throughput evaluation (Table 1). Architecturally, OpenRT functions as a composable toolkit that explicitly decouples core components: models, datasets, attacks, judges, and evaluators under a central orchestrator. The framework integrates 37 attack implementations, covering a broad spectrum of threat models. In the black-box settings, OpenRT supports diverse methodologies, ranging from direct single-turn prompting liu2023autodan; liu2024flipattack to sophisticated multi-turn conversational jailbreaks (e.g., PAIR chao2023jailbreaking, RedQueen jiang2024red, Crescendo russinovich2025great, RACE ying2025reasoning), code-oriented exploitation (e.g., CodeAttack ren2024codeattack), and population-based optimization (e.g., genetic algorithms, GPTFuzzer yu2023gptfuzzer). It also leverages multi-agent and diversity-driven approaches, such as X-Teaming rahman2025x, Rainbow Teaming samvelyan2024rainbow, and EvoSynth chen2025evolve, to maximize the exploration of jailbreak trajectories. In the white-box settings, the framework facilitates gradient-based attacks, including GCG zou2023universal, visual perturbations qi2024visual, and imperceptible jailbreaking gao2025imperceptible. Powered by an asynchronous engine, it unifies API and local model interfaces for scalability. Furthermore, it employs a hybrid evaluation suite that combines rule-based filters with LLM judges, ensuring efficient and robust assessment across diverse safety domains.
@@ 2 Related Work
Red Teaming.  Early approaches to MLLMs safety focused on manual red teaming, where human experts induce harmful outputs through targeted inputs, a process known as jailbreaking (perez2022red; liu2023jailbreaking; weidinger2023sociotechnical; ma2025safety). While effective in uncovering subtle vulnerabilities (Li2024LLMDA; pliny2024), manual methods are limited by scalability, cost, and coverage (bai2022constitutional; ganguli2022red). To address these limitations, automated red teaming has gained attention (yu2023gptfuzzer; mazeika2024harmbench), with early techniques focusing on input space exploration, such as genetic algorithms (liu2023jailbreaking; Lapid2023OpenSU), token-level combinatorial methods (zang2020word), gradient-based optimization (zou2023universal; chen2024llm; geisler2024attacking; wang2025freezevla), and LLM-driven refinement schemes that iteratively improve attack prompts (chao2023jailbreaking; mehrotra2023tree; yu2024llmvirus; zhou2024easyjailbreak; xiao2024tastle). However, these approaches primarily treat jailbreak discovery as search in the input space and remain confined to prompt refinement. Recent work has shifted toward agent-based frameworks that automate not only prompt generation but entire attack strategies, including systems like RedAgent (Xu2024RedAgentRT), ALI-Agent (Zheng2024ALIAgentAL), WildTeaming (jiang2024wildteaming), AutoRedTeamer (zhou2025autoredteamer), AutoDAN-Turbo (liu2024autodan), H4RM3L (doumbouya2024h4rm3l), X-Teaming rahman2025x, and EvoSynth (chen2025evolve), which leverage multi-agent coordination and evolutionary techniques to generate novel attack vectors. Additionally, programmatic attacks such as CodeAttack (jha2023codeattack), which treat code snippets purely as textual input, and co-evolutionary training frameworks like Evo-MARL (pan2025evo) or RL-based adversarial sample generation (zou2019reinforced) have further expanded the range of attack methods. These advancements significantly broaden the scope of automated red teaming, enabling more dynamic and scalable adversarial testing.
@@
Evaluation Frameworks and Benchmarks.  Beyond developing individual attack algorithms, a significant body of work chen2025teleai has focused on creating standardized toolboxes and benchmarks for jailbreak evaluation. EasyJailbreak (zhou2024easyjailbreak), for instance, offers a modular pipeline implementing numerous attack families. To further systematize evaluation, JailbreakBench (chao2024jailbreakbench) and HarmBench (mazeika2024harmbench) provide large-scale test suites and principled metrics for benchmarking adversarial defenses. Specialized frameworks have also emerged, including JailTrickBench (xu2024bag), which focuses on specific jailbreak implementation techniques, and OmniSafeBench-MM (jia2025omnisafebench) for evaluating adversarial robustness in multimodal models. Frameworks like DeepTeam (deepteam) support automated adversarial testing across various attack strategies. In contrast, our work, OpenRT, is distinguished by its comprehensiveness and scale. It uniquely offers a unified and module framework that natively supports diverse MLLMs. By integrating 37 configurable and extensible attack strategies, OpenRT offers a more robust and scalable solution for assessing adversarial vulnerabilities across diverse MLLMs.
@@
EasyJailbreak zhou2024easyjailbreak 11 Text Single $\times$ Low Medium
JailbreakBench chao2024jailbreakbench 4 Text Single $\times$ Medium Low
HarmBench mazeika2024harmbench 18 Text Single $\times$ Medium Medium
JailTrickBench xu2024bag 7 Text Single $\times$ Low Medium
GA GA 6 Text Single & Multi $\times$ Medium Medium
PyRIT munoz2024pyrit 9 Text Single & Multi $\times$ Medium Medium
DeepTeam deepteam 19 Text Single & Multi $\times$ Medium Medium
TeleAI-Safety chen2025teleai 19 Text Single & Multi $\times$ Medium Medium
OmniSafeBench-MM jia2025omnisafebench 13 Image Single $\times$ Medium Medium
@@ 3.1 Preliminaries
where $t_{n+1:n+m}$ represents the target harmful content, $(v^{\prime},\widetilde{t}_{1:n})$ is the adversarial example, $\epsilon$ denotes the perturbation budget for the visual modality madry2017towards, and $\beta$ represents the perplexity threshold for the textual prompt to ensure stealthiness. For text-based jailbreaking, the visual terms $(v,\epsilon)$ are omitted, reducing the optimization solely to the discrete prompt $\widetilde{t}_{1:n}$.
@@ 3.2.1 Target Model
The framework supports two implementations: (1) API-based models compatible with OpenAI-style endpoints, featuring conversation history management, multi-modal input support, and automatic retry mechanisms; (2) Local models with full gradient access for white-box attacks such as GCG (zou2023universal). Multi-turn jailbreaking is enabled through conversation history $\mathcal{H}$, where $r_{k}=\mathcal{Q}(q_{k}\mid\mathcal{H}_{<k})$, controlled by the maintain_history parameter.
@@
GCG zou2023universal 2023 Text Single $\times$ Gradient Optimization
Visual Jailbreak qi2024visual 2023 Image Single $\times$ Gradient Optimization
@@
AutoDAN liu2023autodan 2023 Text Single $\times$ Genetic Algorithm
GPTFuzzer yu2023gptfuzzer 2023 Text Single $\times$ Fuzzing / Mutation
TreeAttack mehrotra2024tree 2023 Text Single $\times$ Tree-Search Optimization
SeqAR yang2025seqar 2024 Text Single $\times$ Genetic Algorithm
RACE ying2025reasoning 2025 Text Single $\times$ Gradient/Genetic Optimization
AutoDAN-R liu2025autodan 2025 Text Single $\times$ Test-Time Scaling
@@
PAIR chao2023jailbreaking 2023 Text Single $\times$ Iterative LLM Optimization
ReNeLLM ding2024wolf 2023 Text Single $\times$ Rewrite & Nesting
DrAttack li2024drattack 2024 Text Single $\times$ Prompt Decomposition
AutoDAN-Turbo liu2024autodan 2024 Text Single $\times$ Genetic + Gradient Guide
@@
CipherChat yuan2023gpt 2023 Text Single $\times$ Cipher/Encryption
CodeAttack ren2024codeattack 2022 Text Single $\times$ Code Encapsulation
Multilingual deng2023multilingual 2023 Text Single $\times$ Low-Resource Language
Jailbroken wei2023jailbroken 2023 Text Single $\times$ Template Combination
ICA wei2023jailbreak 2023 Text Single $\times$ In-Context Demonstration
FlipAttack liu2024flipattack 2024 Text Single $\times$ Token Flipping / Masking
Mousetrap yao2025mousetrap 2025 Text Single $\times$ Logic Nesting / Obfuscation
Prefill li2025prefill 2025 Text Single $\times$ Prefix Injection
@@
DeepInception li2023deepinception 2023 Text Single $\times$ Hypnosis or Nested Scene
Crescendo russinovich2025great 2024 Text Multi $\times$ Multi-turn Steering
RedQueen jiang2024red 2024 Text Multi $\times$ Concealed Knowledge
CoA yang2024chain 2024 Text Multi $\times$ Chain of Attack
@@
FigStep gong2025figstep 2023 Image Single $\times$ Typography / OCR
QueryRelevant liu2023query 2024 Image Single $\times$ Visual Prompt Injection
IDEATOR wang2025ideator 2024 Image Single $\times$ Visual Semantics
MML wang2025jailbreak 2024 Image Single $\times$ Cross‑Modal Encryption
HADES li2024images 2024 Image Single $\times$ Visual Vulnerability Amplification
HIMRD ma2025heuristic 2024 Image Single $\times$ Multi-Modal Risk Distribution
JOOD jeong2025playing 2025 Image Single $\times$ OOD Transformation
SI zhao2025jailbreaking 2025 Image Single $\times$ Shuffle Inconsistency Optimization
CS-DJ yang2025distraction 2025 Image Single $\times$ Multi‑Level Visual Distraction
@@
ActorAttack ren2024derail 2024 Text Multi $\checkmark$ Actor-Based Steering
Rainbow Teaming samvelyan2024rainbow 2024 Text Multi $\checkmark$ Diversity-Driven Search
X-Teaming rahman2025x 2025 Text Multi $\checkmark$ Cooperative Exploration
EvoSynth chen2025evolve 2025 Text Multi $\checkmark$ Code-Level Evolutionary Synthesis
@@ 3.2.2 Dataset
There are different types of datasets tailored to specific use cases. The first type, the StaticDataset, is an in-memory dataset suitable for small-scale testing. On the other hand, the JSONLDataset supports streaming loading of large-scale evaluation benchmarks (e.g., AdvBench (zou2023universal), HarmBench (mazeika2024harmbench)).

### v2

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) Team (2025c); OpenAI (2025b); Anthropic (2025); Google (2025); Comanici et al. (2025); AI (2025a); xAI (2025) are increasingly powering real-world applications, including conversational assistants Akhoroz and Yildirim (2025); Fischer et al. (2024), code copilots Ross et al. (2023), and search agents Gong et al. (2024); Xi et al. (2025). To mitigate harmful behavior, these systems are often equipped with safety alignment mechanisms Lab (2025); Qi et al. (2025); Wang et al. (2025d) and safeguard policies Inan et al. (2023); Wang et al. (2024b, 2025b); Zhao et al. (2025a). Despite being widely adopted, conventional defenses such as system prompts Sharma et al. (2025), safety filters Huang et al. (2024); Gong et al. (2025a), and refusal-aware fine-tuning Li et al. (2024a); Zhang et al. (2025), remain vulnerable to adversarial attacks Chen et al. (2025b); Rahman et al. (2025); Wang et al. (2025e), revealing a significant gap between perceived safety and empirical worst-case vulnerabilities.
@@
Despite significant advances in jailbreak techniques Qi et al. (2024); Zou et al. (2023); Chao et al. (2025), the ecosystem for systematically evaluating adversarial robustness remains fragmented. Most existing red-teaming frameworks Zhou et al. (2024); Chao et al. (2024); Mazeika et al. (2024); Xu et al. (2024b); Jia et al. (2025) focus on a narrow subset of attacks, limited threat models, or a small selection of target models. As the number and variety of red-teaming approaches grow, including evolutionary strategies Chen et al. (2025b), multi-modal jailbreaks Wang et al. (2025a); Gong et al. (2025b), multi-turn optimization Samvelyan et al. (2024); Liu et al. (2025b), and multi-agent coordination Rahman et al. (2025); Ren et al. (2024b), the lack of a unified experimental framework has become a critical bottleneck. Such fragmentation undermines reproducible benchmarking and limits the systematic evaluation of vulnerabilities across models. As a result, it remains difficult to establish standardized baselines for attack efficacy or quantify the consistency of safety failures across diverse models.
@@
In this work, we introduce OpenRT, a modular and extensible framework designed for the red teaming of MLLMs. Unlike existing toolboxes often limited to a narrow subset of classic attacks, OpenRT supports massively parallel jailbreaking in both white-box and black-box settings, engineered specifically for high-throughput evaluation (Table 1). Architecturally, OpenRT functions as a composable toolkit that explicitly decouples core components: models, datasets, attacks, judges, and evaluators under a central orchestrator. The framework integrates 37 attack implementations, covering a broad spectrum of threat models. In the black-box settings, OpenRT supports diverse methodologies, ranging from direct single-turn prompting Liu et al. (2024a, b) to sophisticated multi-turn conversational jailbreaks (e.g., PAIR Chao et al. (2025), RedQueen Jiang et al. (2024b), Crescendo Russinovich et al. (2025), RACE Ying et al. (2025)), code-oriented exploitation (e.g., CodeAttack Ren et al. (2024a)), and population-based optimization (e.g., genetic algorithms, GPTFuzzer Yu et al. (2023)). It also leverages multi-agent and diversity-driven approaches, such as X-Teaming Rahman et al. (2025), Rainbow Teaming Samvelyan et al. (2024), and EvoSynth Chen et al. (2025b), to maximize the exploration of jailbreak trajectories. In the white-box settings, the framework facilitates gradient-based attacks, including GCG Zou et al. (2023), visual perturbations Qi et al. (2024), and imperceptible jailbreaking Gao et al. (2025). Powered by an asynchronous engine, it unifies API and local model interfaces for scalability. Furthermore, it employs a hybrid evaluation suite that combines rule-based filters with LLM judges, ensuring efficient and robust assessment across diverse safety domains.
@@ 2 Related Work
Red Teaming.  Early approaches to MLLMs safety focused on manual red teaming, where human experts induce harmful outputs through targeted inputs, a process known as jailbreaking (Perez et al., 2022; Liu et al., 2023b; Weidinger et al., 2023; Ma et al., 2025b). While effective in uncovering subtle vulnerabilities (Li et al., 2024b; the Prompter, 2024), manual methods are limited by scalability, cost, and coverage (Bai et al., 2022; Ganguli et al., 2022). To address these limitations, automated red teaming has gained attention (Yu et al., 2023; Mazeika et al., 2024), with early techniques focusing on input space exploration, such as genetic algorithms (Liu et al., 2023b; Lapid et al., 2024), token-level combinatorial methods (Zang et al., 2020), gradient-based optimization (Zou et al., 2023; Chen et al., 2024; Geisler et al., 2024; Wang et al., 2025c), and LLM-driven refinement schemes that iteratively improve attack prompts (Chao et al., 2025; Mehrotra et al., 2023; Yu et al., 2024; Zhou et al., 2024; Xiao et al., 2024). However, these approaches primarily treat jailbreak discovery as search in the input space and remain confined to prompt refinement. Recent work has shifted toward agent-based frameworks that automate not only prompt generation but entire attack strategies, including systems like RedAgent (Xu et al., 2024a), ALI-Agent (Wang et al., 2024a), WildTeaming (Jiang et al., 2024a), AutoRedTeamer (Zhou et al., 2025), AutoDAN-Turbo (Liu et al., 2025b), H4RM3L (Doumbouya et al., 2024), X-Teaming Rahman et al. (2025), and EvoSynth (Chen et al., 2025b), which leverage multi-agent coordination and evolutionary techniques to generate novel attack vectors. Additionally, programmatic attacks such as CodeAttack (Jha and Reddy, 2023), which treat code snippets purely as textual input, and co-evolutionary training frameworks like Evo-MARL (Pan et al., 2025) or RL-based adversarial sample generation (Zou et al., 2019) have further expanded the range of attack methods. These advancements significantly broaden the scope of automated red teaming, enabling more dynamic and scalable adversarial testing.
@@
Evaluation Frameworks and Benchmarks.  Beyond developing individual attack algorithms, a significant body of work Chen et al. (2025a) has focused on creating standardized toolboxes and benchmarks for jailbreak evaluation. EasyJailbreak (Zhou et al., 2024), for instance, offers a modular pipeline implementing numerous attack families. To further systematize evaluation, JailbreakBench (Chao et al., 2024) and HarmBench (Mazeika et al., 2024) provide large-scale test suites and principled metrics for benchmarking adversarial defenses. Specialized frameworks have also emerged, including JailTrickBench (Xu et al., 2024b), which focuses on specific jailbreak implementation techniques, and OmniSafeBench-MM (Jia et al., 2025) for evaluating adversarial robustness in multimodal models. Frameworks like DeepTeam (deepteam, 2025) support automated adversarial testing across various attack strategies. In contrast, our work, OpenRT, is distinguished by its comprehensiveness and scale. It uniquely offers a unified and module framework that natively supports diverse MLLMs. By integrating 37 configurable and extensible attack strategies, OpenRT offers a more robust and scalable solution for assessing adversarial vulnerabilities across diverse MLLMs.
@@
EasyJailbreak Zhou et al. (2024) 11 Text Single $\times$ Low Medium
JailbreakBench Chao et al. (2024) 4 Text Single $\times$ Medium Low
HarmBench Mazeika et al. (2024) 18 Text Single $\times$ Medium Medium
JailTrickBench Xu et al. (2024b) 7 Text Single $\times$ Low Medium
GA GeneralAnalysis (2025) 6 Text Single & Multi $\times$ Medium Medium
PyRIT Munoz et al. (2024) 9 Text Single & Multi $\times$ Medium Medium
DeepTeam deepteam (2025) 19 Text Single & Multi $\times$ Medium Medium
TeleAI-Safety Chen et al. (2025a) 19 Text Single & Multi $\times$ Medium Medium
OmniSafeBench-MM Jia et al. (2025) 13 Image Single $\times$ Medium Medium
@@ 3.1 Preliminaries
where $t_{n+1:n+m}$ represents the target harmful content, $(v^{\prime},\widetilde{t}_{1:n})$ is the adversarial example, $\epsilon$ denotes the perturbation budget for the visual modality Madry et al. (2017), and $\beta$ represents the perplexity threshold for the textual prompt to ensure stealthiness. For text-based jailbreaking, the visual terms $(v,\epsilon)$ are omitted, reducing the optimization solely to the discrete prompt $\widetilde{t}_{1:n}$.
@@ 3.2.1 Target Model
The framework supports two implementations: (1) API-based models compatible with OpenAI-style endpoints, featuring conversation history management, multi-modal input support, and automatic retry mechanisms; (2) Local models with full gradient access for white-box attacks such as GCG (Zou et al., 2023). Multi-turn jailbreaking is enabled through conversation history $\mathcal{H}$, where $r_{k}=\mathcal{Q}(q_{k}\mid\mathcal{H}_{<k})$, controlled by the maintain_history parameter.
@@
GCG Zou et al. (2023) 2023 Text Single $\times$ Gradient Optimization
Visual Jailbreak Qi et al. (2024) 2023 Image Single $\times$ Gradient Optimization
@@
AutoDAN Liu et al. (2024a) 2023 Text Single $\times$ Genetic Algorithm
GPTFuzzer Yu et al. (2023) 2023 Text Single $\times$ Fuzzing / Mutation
TreeAttack Mehrotra et al. (2024) 2023 Text Single $\times$ Tree-Search Optimization
SeqAR Yang et al. (2025a) 2024 Text Single $\times$ Genetic Algorithm
RACE Ying et al. (2025) 2025 Text Single $\times$ Gradient/Genetic Optimization
AutoDAN-R Liu and Xiao (2025) 2025 Text Single $\times$ Test-Time Scaling
@@
PAIR Chao et al. (2025) 2023 Text Single $\times$ Iterative LLM Optimization
ReNeLLM Ding et al. (2024) 2023 Text Single $\times$ Rewrite & Nesting
DrAttack Li et al. (2024c) 2024 Text Single $\times$ Prompt Decomposition
AutoDAN-Turbo Liu et al. (2025b) 2024 Text Single $\times$ Genetic + Gradient Guide
@@
CipherChat Yuan et al. (2024) 2023 Text Single $\times$ Cipher/Encryption
CodeAttack Ren et al. (2024a) 2022 Text Single $\times$ Code Encapsulation
Multilingual Deng et al. (2023) 2023 Text Single $\times$ Low-Resource Language
Jailbroken Wei et al. (2023a) 2023 Text Single $\times$ Template Combination
ICA Wei et al. (2023b) 2023 Text Single $\times$ In-Context Demonstration
FlipAttack Liu et al. (2024b) 2024 Text Single $\times$ Token Flipping / Masking
Mousetrap Yao et al. (2025) 2025 Text Single $\times$ Logic Nesting / Obfuscation
Prefill Li et al. (2025) 2025 Text Single $\times$ Prefix Injection
@@
DeepInception Li et al. (2023) 2023 Text Single $\times$ Hypnosis or Nested Scene
Crescendo Russinovich et al. (2025) 2024 Text Multi $\times$ Multi-turn Steering
RedQueen Jiang et al. (2024b) 2024 Text Multi $\times$ Concealed Knowledge
CoA Yang et al. (2024) 2024 Text Multi $\times$ Chain of Attack
@@
FigStep Gong et al. (2025b) 2023 Image Single $\times$ Typography / OCR
QueryRelevant Liu et al. (2023a) 2024 Image Single $\times$ Visual Prompt Injection
IDEATOR Wang et al. (2025a) 2024 Image Single $\times$ Visual Semantics
MML Wang et al. (2025f) 2024 Image Single $\times$ Cross‑Modal Encryption
HADES Li et al. (2024d) 2024 Image Single $\times$ Visual Vulnerability Amplification
HIMRD Ma et al. (2025a) 2024 Image Single $\times$ Multi-Modal Risk Distribution
JOOD Jeong et al. (2025) 2025 Image Single $\times$ OOD Transformation
SI Zhao et al. (2025b) 2025 Image Single $\times$ Shuffle Inconsistency Optimization
CS-DJ Yang et al. (2025b) 2025 Image Single $\times$ Multi‑Level Visual Distraction
@@
ActorAttack Ren et al. (2024b) 2024 Text Multi $\checkmark$ Actor-Based Steering
Rainbow Teaming Samvelyan et al. (2024) 2024 Text Multi $\checkmark$ Diversity-Driven Search
X-Teaming Rahman et al. (2025) 2025 Text Multi $\checkmark$ Cooperative Exploration
EvoSynth Chen et al. (2025b) 2025 Text Multi $\checkmark$ Code-Level Evolutionary Synthesis
@@ 3.2.2 Dataset
There are different types of datasets tailored to specific use cases. The first type, the StaticDataset, is an in-memory dataset suitable for small-scale testing. On the other hand, the JSONLDataset supports streaming loading of large-scale evaluation benchmarks (e.g., AdvBench (Zou et al., 2023), HarmBench (Mazeika et al., 2024)).